Starkey (STAR1)
Site Details
Country:United States
State:Oregon (pop. 2842321)
County:Union County (pop. 23598)
Latitude:45.2249
Longitude:-118.5129
Elevation:1259m above MSL
EPA Code:410610010
Agency:US Forest Service
Sponsor:USFS
Region:Hells Canyon
Land Use:Unknown
AQCR:None
Dates:3/2000 - Present
SiteID:76
SiteCode:STAR1
SiteName:Starkey
Maps:Google, MSN, Yahoo
Site History
10/30/2003:Prescribed fire continues 1 mile N and 1 mile W
10/11/2002:Prescribed fire in area for three weeks
10/06/2003:Prescribed fire, increased traffic due to hunt season
10/01/2003:A1 Module Flow Calibration
10/01/2003:B1 Module Flow Calibration
10/01/2003:C1 Module Flow Calibration
10/01/2003:C1 Module pump replaced
10/01/2003:D1 Module Flow Calibration
10/01/2003:D1 Module pump replaced
10/01/2003:Photos taken
10/01/2003:Routine annual maintenance
09/27/2000:logging activity this week
09/23/2002:Prescribed fire in area this week
09/08/2001:B1 Module Flow Calibration
09/08/2001:C1 Module Flow Calibration
09/08/2001:D1 Module Flow Calibration
09/08/2001:Routine annual maintenance
09/08/2001:A1 Module Flow Calibration
09/06/2003:Smoke from fire outside immediate area blown over this past week
08/31/2003:bow hunting started, increased activity past collection site
08/31/2000:A1 Module Flow Calibration
08/31/2000:B1 Module Flow Calibration
08/31/2000:C1 Module Flow Calibration
08/31/2000:D1 Module Flow Calibration
08/26/2001:Continued traffic from log and chip trucks
08/23/2001:C1 Module pump replaced
08/22/2003:Fire north and west, but none in immediate area
08/19/2000:Smoke from distant fire
08/17/2001:Increased road traffic
08/14/2001:Fire 30 miles west of the site
08/06/2002:hazy from area fire, increased traffic this week
08/02/2001:Road traffic from chip trucks continues
07/24/2001:A1 Module electronics box replaced
07/20/2004:A1 Module Flow Calibration
07/20/2004:B1 Module Flow Calibration
07/20/2004:C1 Module Flow Calibration
07/20/2004:D1 Module Flow Calibration
07/20/2004:Photos taken
07/20/2004:Routine annual maintenance
07/12/2001:Road traffic from chip trucks continues for two weeks
07/10/2002:Hazy from fire
06/24/2002:A1 Module Flow Calibration
06/24/2002:A1 Module pump replaced
06/24/2002:B1 Module Flow Calibration
06/24/2002:B1 Module pump replaced
06/24/2002:C1 Module Flow Calibration
06/24/2002:D1 Module Flow Calibration
06/24/2002:Photos taken
06/24/2002:Routine annual maintenance
06/24/2002:Temperature probe replaced
06/20/2003:Log truck traffic continues until 08/28/2003
06/17/2004:A1 Module area increased from 2.20 to 3.53 sq cm
05/10/2001:Prescribed fire
04/05/2000:Prescribed fire 50 yards north
03/22/2000:Prescribed fire 50 yards north
03/18/2000:Prescribed fire 50 yards north
03/12/2000:Version 2 sampler installed
03/07/2000:New IMPROVE site, version 2 sampler installed
03/07/2000:A1 Module Flow Calibration
03/07/2000:B1 Module Flow Calibration
03/07/2000:C1 Module Flow Calibration
03/07/2000:D1 Module Flow Calibration
03/07/2000:Photos taken